### v1

@@ 3.4 Effects of local vessel stiffness and flow path
3.5 Effects of stent grafting and bypass angle
The effects of the local vessel stiffness (Young’s modulus) and stent-grafted length on the phase difference in the L/RVA were further investigated using the postoperative vascular tree model. Simulations were performed for different Young’s modulus valuses $E/E_{\mathrm{s}}$ ($=0.01,0.1,1,10$) in the stent-grafted artery with standard length $L_{0}=40$ mm, where $E_{\mathrm{s}}$ ($=10$ MPa) was the original Young’s modulus in the stent-grafted region. The ratio of the phase difference $\delta/T$ remained almost the same (less than $0.2$%) in both the L/RVA. Even when the length of the stent-grafted region increased to $L/L_{0}=1,2.25,3.5,4.75,6$, the results of $\delta/T$ remained the same (less than $0.1$%; data not shown), where $L_{0}$ ($=40$ mm) was the original length of the stent-grafted region.
The phase delay in the LVA was also insensitive to the bypass angle $\theta$, defined at the branch point between the vessel and bypass, especially for $\theta\geq 60$ deg as shown in Figure 7. Note that the relative difference of the mean flow rate in the RVA obtained with the reference angle ($30$ deg) and that obtained with $120$ deg at the branch angle of the bypass was smaller than 1% (data not shown). The results were obtained with the postoperative vascular tree model with a standard stent length of $L_{0}=40$ mm and a Young’s modulus of $E_{\mathrm{s}}=10$ MPa. These results, including those shown in Figure 6(C), suggest that the phase delay of the flow rate in the VA between before and after surgery (Figure 3) arises mostly from the alteration of the flow path, i.e., by the new bypass and embolization rather than by local vessel stiffness due to stent grafting.
Note that the effect of the stiffness (or Young’s modulus $E$) of the straight tube on the fluid velocity is described in the Appendix, and the results showed that the pulse-wave speed $c$ increased with $E$ with the ratio of velocity $O(c/U)>10^{2}$ (Figure A2D).
Figure 7: Normalized phase difference of the flow rate $\delta/T$ in both the LVA and RVA between the preoperative and postoperative vascular tree models with different branch angles $\theta$ at the vessel-bypass.
@@ 4 Discussion
A previous clinical study by Goto et al. (2019) showed hemodynamic changes in blood flow through the L/RVA after $1$dTEVAR, while postoperative total intracranial blood flow was almost the same as that measured preoperatively. This might be due to systemic hemodynamic compensation causing cerebral blood flow to be strictly maintained through cerebral autoregulation (Seifert and Secher, 2011). Despite the maintained flow mass in patients after $1$dTEVAR, experimental measurements demonstrated phase differences of the flow rate in the LVA but not in the RVA between pre- and postoperative patients undergoing $1$dTEVAR (Figure 3). It is expected that the aforementioned hemodynamic compensation and the phase differences in the VA in postoperative patients might be due to two mechanical factors: structural alterations in flow paths due to the new bypass and embolization, and an increase of local vessel stiffness due to stent grafting. However, much is still unknown about this matter, in particular about the relative contribution of these two factors to the phase of the flow rate in the VA between pre- and postoperative patients undergoing $1$dTEVAR. To explore this issue, A $1$D model was used to numerically investigate flow rates in the VA, both before and after $1$dTEVAR. Using a postoperative arterial tree model (Figure 1(B)), The numerical results demonstrated a phase delay in the LVA compared to preoperative models, as shown in Figure 5(A). Stent grafting, which is characterized by a locally increasing Young’s modulus, contributed negligibly to the phase delay in the LVA (see §3.4). Thus, the experimentally observed phase delay (Figure 3(A)) was mainly caused by alteration of the flow path, i.e., by the new bypass and embolization, rather by local vessel stiffness due to stent grafting (Figure 6).
@@
In this simulation, branch angles were uniformly set to $30$ deg for simplicity. A previous experimental study showed that the energy loss at bifurcations was very small as reported in (Matthys et al., 2007). This is because the branch angle only affects the total pressure continuity (equations 18(b, c) and 20(b, c), see the Appendix), and the order of magnitude of the static pressure term $p$ is much greater than those of the other terms. Indeed, the relative difference of the mean flow rate in the RVA obtained with the reference angle ($30$ deg) and that obtained with $120$ deg at the branch angle of the bypass is smaller than 1% (data not shown). Thus, at least in this model, the phase delay in the LVA was insensitive to the bypass angle $\theta$. We did not consider postoperative shape deformation and diameter changes due to device placement. In the future we will perform systematic analyses of the effect of postoperative vascular configurations on the phase change.
@@
In this study, the Young’s modulus in model vasculatures was also approximated using equation (4). Referring to previous numerical assessments of the mechanical behavior of ePTEF, we set the same order of magnitude of the Young’s modulus for the stent-grafted area (i.e., $E_{\mathrm{s}}=10$ MPa), which was approximately 100 times larger than that of vessels with similar diameter. Considering different Young’s moduli of PET stent from $O(10^{0})$ MPa (Kan et al., 2021; Kleinstreuer et al., 2008; Ramella et al., 2022) to $O(10^{3})$ MPa (Ma et al., 2018; Ramella et al., 2022) in previous numerical studies, the simulations were performed for different orders of magnitude of Young’s modulus in an ePTFE-stent-grafted area ($E/E_{\mathrm{s}}=0.01-10$). However, the ratio of the phase difference $\delta/T$ remained almost the same (less than 0.2%) in both the L/RVA. Since model parameters in the bifurcation model ($\gamma_{1}$ and $\gamma_{2}$ in equations 18(b, c) and 20(b, c), see the Appendix) and in Young’s modulus ($k_{1},k_{2}$, and $k_{3}$ in equation (4)) were fixed in this study, further precise analysis of the effects of bifurcation angles and Young’s modulus of the stent should be performed experimentally to confirm whether those mechanical factors have less impact on the phase differences.
@@
In this simulation, we simply modeled arterial trees from the heart (or LV) to eight end terminals, whose diameters and lengths were determined so that the time average flow rate became similar to that obtained in experimental measurements by Bogren and Buonocore (1994) or Goto et al. (2019) (Figure 4). Even with the simplest boundary condition at the end terminals (no-reflection condition, i.e., $W_{-}=0$), our numerical model reproduced the phase delay of the flow rate in the LVA without the phase difference in the RVA in postoperative patients, and also clarified that the phase delay arose mostly from alteration of the flow path (Figure 6). Further precise boundary conditions for both the inlet (heart) and outlets will leads to patient-specific analyses preoperatively and to the evaluation of hemodynamics postoperatively, topics that we will address in the future.
@@ 5 Conclusion
A 1D model was used to numerically investigate blood flow rates, with the goal of explaining experimental evidence on the phase delay of the flow rate in the LVA but not in the RVA after $1$dTEVAR relative to before. The numerical model can reproduce the flow distribution in the major arteries from the heart, and can capture the flow rate ratio in the L/RVA in both pre- and postoperative patients. The numerical results showed that the phase delay was mainly caused by the bypass, i.e., by alteration of the flow path, rather than by stent grafting, i.e., the change of local vessel stiffness. Bypass angles and the effects of the length and Young’s modulus of the stent were also investigated, but all were insensitive to the phase delay. We hope that our numerical results will provide fundamental knowledge about therapeutic decisions for dTEVAR.
@@ Acknowledgements
A1. Methodology
@@
Flow conditions along the lines ($1\to 2$) (see Figure A1(A)) are described by mass conservation and total pressure continuity (Formaggia et al., 2003):
@@
and $W_{1}^{n+1}(L)$ and $W_{2}^{n+1}(0)$ are derived from values at the previous time step at the distal end (denoted by $L$) and proximal end (denoted by 0) of an artery by extrapolating the outgoing Riemann invariants along the characteristic lines (Formaggia et al., 2003):
@@
Flow conditions at the bifurcations ($3\to 1+2$) (Figure A1(B)) are described as (Formaggia et al., 2003):
@@
A2. Weakly non-linear form
@@
Equation (23) predicts that wave propagations, including reflective wave $W_{-}$, are negligible for the small flow rate $dQ\approx 0$ and small vessel contraction $A\approx A_{0}$. Indeed, the maximum values of $|W_{\pm}|$ in the middle point of the straight tube almost collapse on those obtained with $p_{a}=0$ when the maximum inlet flow rate $Q_{\mathrm{in}}^{\mathrm{max}}$ decreases (Figure A2(B)), where the tube length and reference radius are set to $L=1$ m and $r_{0}=10^{-3}$ m, respectively, and the Young’s modulus is $E_{0}=0.1$ MPa, referring to the value of aortic elasticity in conscious dogs (Armentano et al., 1991). The inlet pressure wave form (7) and no-reflection condition for the outlet $W_{-}=0$ are also considered. The small maximum inlet inflow rate $Q_{\mathrm{in}}^{\mathrm{max}}$ is controlled by the inlet pressure amplitude $p_{a}$ (see Figure A2(A)).

### v2

@@ 4 Discussion
A previous clinical study by Goto et al. [3] showed hemodynamic changes in blood flow through the L/RVA after $1$dTEVAR, while postoperative total intracranial blood flow was almost the same as that measured preoperatively. This might be due to systemic hemodynamic compensation causing cerebral blood flow to be strictly maintained through cerebral autoregulation [44]. Despite the maintained flow mass in patients after $1$dTEVAR, experimental measurements demonstrated phase differences of the flow rate in the LVA but not in the RVA between pre- and postoperative patients undergoing $1$dTEVAR (Figure 3). To the best of our knowledge, this study is the first report to describe phase differences of the flow rate in LVAs after 1dTEVAR with AxA-AxA bypass. It is expected that the aforementioned hemodynamic compensation and the phase differences in the VA in postoperative patients might be due to two mechanical factors: structural alterations in flow paths due to the new bypass and embolization, and an increase of local vessel stiffness due to stent grafting. However, much is still unknown about this matter, in particular about the relative contribution of these two factors to the phase of the flow rate in the VA between pre- and postoperative patients undergoing $1$dTEVAR. To explore this issue, A $1$D model was used to numerically investigate flow rates in the VA, both before and after $1$dTEVAR. Using a postoperative arterial tree model (Figure 1(B)), The numerical results demonstrated a phase delay in the LVA compared to preoperative models, as shown in Figure 5(A). Stent grafting, which is characterized by a locally increasing Young’s modulus, contributed negligibly to the phase delay in the LVA (see §3.4). Thus, the experimentally observed phase delay (Figure 3(A)) was mainly caused by alteration of the flow path, i.e., by the new bypass and embolization, rather by local vessel stiffness due to stent grafting (Figure 6).
As reported in previous study [3], the 30-day mortality and inpatient death rates were both 0% in patients after $1$dTEVAR. Furthermore, no postoperative symptomatic stroke was observed in postoperative patients. Thus, a phase delay in the flow rate of the LVA after 1dTEVAR, relative to before, does not affect the postoperative stroke, at least in the short-term postoperative prognosis. Although further statistical analysis is required due to the small number of enrolled patients, at the present, there is no clinical evidence that postoperative phase delay may be critical to their quality of life (QOL). A follow-up study is needed to clarify whether the postoperative phase delay in LVA can be related to long-term cerebral complications or prognosis. To evaluate the long-term impact of such a significant phase difference on posterior cerebral circulation, follow-up MRI should be performed periodically, which is however our future study.
In general, the morphometric properties of the aorta differ along its length. In healthy humans, helical computed tomography has showed that the maximal aortic diameter is located in the ascending aorta, just distal to the aortic valve sinus and proximal to the innominate artery [49]. The aortic diameter then progressively decreases along the thoracic aorta and continues to decrease from the infrarenal abdominal aorta to the lower abdominal aorta [49, 50]. Similarly, the overall thickness of the aortic wall also decreases along the thoracic aorta but remains constant in the abdominal aorta [51]. These findings are qualitatively similar to our measurements of reference radius $r_{0}$ in Table 2 and consistent with our modeling (equations(7) and (8)). It is also known that aging causes morphometric changes in the aforementioned geometric parameters of the aorta (diameter, length, and thickness) [52]. Additionally, aortic stiffness increases significantly with age and is associated with an elevated risk for various adverse clinical outcomes, e.g., heart disease, dementia, and kidney disease [53]. In this study, the geometrical parameters in Table 2 and mechanical property represented by Young’s modulus referred to normal subjects. Given that preoperative patients have the aortic aneurysm and have potential complications, selecting these parameters for preoperative patients will lead to a more rigorous estimation of the phase delay. However, a systematic parametric study of these parameters is beyond the scope of the present work.
@@
In this simulation, branch angles were uniformly set to $30$ deg for simplicity. A previous experimental study showed that the energy loss at bifurcations was very small as reported in [45]. This is because the branch angle only affects the total pressure continuity (equations 18(b, c) and 20(b, c), see the Appendix §A.1), and the order of magnitude of the static pressure term $p$ is much greater than those of the other terms. Indeed, the relative difference of the mean flow rate in the RVA obtained with the reference angle ($30$ deg) and that obtained with $120$ deg at the branch angle of the bypass is smaller than 1% (data not shown). Thus, at least in this model, the phase delay in the LVA was insensitive to the bypass angle $\theta$. We did not consider postoperative shape deformation and diameter changes due to device placement. In the future we will perform systematic analyses of the effect of postoperative vascular configurations on the phase change.
@@
In this study, the Young’s modulus in model vasculatures was also approximated using equation (4). Referring to previous numerical assessments of the mechanical behavior of ePTEF, we set the same order of magnitude of the Young’s modulus for the stent-grafted area (i.e., $E_{\mathrm{s}}=10$ MPa), which was approximately 100 times larger than that of vessels with similar diameter. Considering different Young’s moduli of PET stent from $O(10^{0})$ MPa [42, 46, 47] to $O(10^{3})$ MPa [47, 48] in previous numerical studies, the simulations were performed for different orders of magnitude of Young’s modulus in an ePTFE-stent-grafted area ($E/E_{\mathrm{s}}=0.01-10$). However, the ratio of the phase difference $\delta/T$ remained almost the same (less than 0.2%) in both the L/RVA. Since model parameters in the bifurcation model ($\gamma_{1}$ and $\gamma_{2}$ in equations 18(b, c) and 20(b, c), see the Appendix §A.2) and in Young’s modulus ($k_{1},k_{2}$, and $k_{3}$ in equation (4)) were fixed in this study, further precise analysis of the effects of bifurcation angles and Young’s modulus of the stent should be performed experimentally to confirm whether those mechanical factors have less impact on the phase differences.
@@
In this simulation, we simply modeled arterial trees from the heart (or LV) to eight end terminals, whose diameters and lengths were determined so that the time average flow rate became similar to that obtained in experimental measurements by Bogren & Buonocore [38] or Goto et al. [3] (Figure 4). Even with the simplest boundary condition at the end terminals (no-reflection condition, i.e., $W_{-}=0$), our numerical model reproduced the phase delay of the flow rate in the LVA without the phase difference in the RVA in postoperative patients, and also clarified that the phase delay arose mostly from alteration of the flow path (Figure 6). Further precise boundary conditions for both the inlet (heart) and outlets will leads to patient-specific analyses preoperatively and to the evaluation of hemodynamics postoperatively, topics that we will address in the future.
@@ 5 Conclusion
This study is the first report to describe phase differences of the flow rate in LVAs after 1dTEVAR with AxA-AxA bypass. A 1D model was used to numerically investigate blood flow rates, with the goal of explaining experimental evidence on the phase delay of the flow rate in the LVA but not in the RVA after $1$dTEVAR relative to before. The numerical model can reproduce the flow distribution in the major arteries from the heart, and can capture the flow rate ratio in the L/RVA in both pre- and postoperative patients. The numerical results showed that the phase delay was mainly caused by the bypass, i.e., by alteration of the flow path, rather than by stent grafting, i.e., the change of local vessel stiffness. Bypass angles and the effects of the length and Young’s modulus of the stent were also investigated, but all were insensitive to the phase delay. We hope that our numerical results will provide fundamental knowledge about therapeutic decisions for dTEVAR.
@@ Acknowledgements
A1. Effects of stent grafting and bypass angle
The effects of the local vessel stiffness (Young’s modulus) and stent-grafted length on the phase difference in the L/RVA were investigated using the postoperative vascular tree model. Simulations were performed for different Young’s modulus valuses $E/E_{\mathrm{s}}$ ($=0.01,0.1,1,10$) in the stent-grafted artery with standard length $L_{0}=40$ mm, where $E_{\mathrm{s}}$ ($=10$ MPa) was the original Young’s modulus in the stent-grafted region. The ratio of the phase difference $\delta/T$ remained almost the same (less than $0.2$%) in both the L/RVA. Even when the length of the stent-grafted region increased to $L/L_{0}=1,2.25,3.5,4.75,6$, the results of $\delta/T$ remained the same (less than $0.1$%; data not shown), where $L_{0}$ ($=40$ mm) was the original length of the stent-grafted region.
The phase delay in the LVA was also insensitive to the bypass angle $\theta$, defined at the branch point between the vessel and bypass, especially for $\theta\geq 60$ deg as shown in Figure A1. Note that the relative difference of the mean flow rate in the RVA obtained with the reference angle ($30$ deg) and that obtained with $120$ deg at the branch angle of the bypass was smaller than 1% (data not shown). The results were obtained with the postoperative vascular tree model with a standard stent length of $L_{0}=40$ mm and a Young’s modulus of $E_{\mathrm{s}}=10$ MPa. These results, including those shown in Figure 6(C), suggest that the phase delay of the flow rate in the VA between before and after surgery (Figure 3) arises mostly from the alteration of the flow path, i.e., by the new bypass and embolization rather than by local vessel stiffness due to stent grafting.
Note that the effect of the stiffness (or Young’s modulus $E$) of the straight tube on the fluid velocity is described in §A.3.and the results showed that the pulse-wave speed $c$ increased with $E$ with the ratio of velocity $O(c/U)>10^{2}$ (Figure A3D).
Figure A1: Normalized phase difference of the flow rate $\delta/T$ in both the LVA and RVA between the preoperative and postoperative vascular tree models with different branch angles $\theta$ at the vessel-bypass.
A2. Methodology
@@
Flow conditions along the lines ($1\to 2$) (see Figure A2(A)) are described by mass conservation and total pressure continuity [22]:
@@
and $W_{1}^{n+1}(L)$ and $W_{2}^{n+1}(0)$ are derived from values at the previous time step at the distal end (denoted by $L$) and proximal end (denoted by 0) of an artery by extrapolating the outgoing Riemann invariants along the characteristic lines [22]:
@@
Flow conditions at the bifurcations ($3\to 1+2$) (Figure A2(B)) are described as [22]:
@@
A3. Weakly non-linear form
@@
Equation (23) predicts that wave propagations, including reflective wave $W_{-}$, are negligible for the small flow rate $dQ\approx 0$ and small vessel contraction $A\approx A_{0}$. Indeed, the maximum values of $|W_{\pm}|$ in the middle point of the straight tube almost collapse on those obtained with $p_{a}=0$ when the maximum inlet flow rate $Q_{\mathrm{in}}^{\mathrm{max}}$ decreases (Figure A3(B)), where the tube length and reference radius are set to $L=1$ m and $r_{0}=10^{-3}$ m, respectively, and the Young’s modulus is $E_{0}=0.1$ MPa, referring to the value of aortic elasticity in conscious dogs [41]. The inlet pressure wave form (7) and no-reflection condition for the outlet $W_{-}=0$ are also considered. The small maximum inlet inflow rate $Q_{\mathrm{in}}^{\mathrm{max}}$ is controlled by the inlet pressure amplitude $p_{a}$ (see Figure A3(A)).